The scores of all applicants taking an aptitude test required by a law school have a normal distribution with a mean of 420 and a standard deviation of 100. A random sample of 25 scores is taken.

e. The probability is 0.05 that the sample standard deviation of the scores is higher than what number?

f. The probability is 0.05 that the sample standard deviation of the scores is lower than what number?

g. If a sample of 50 test scores had been taken, would the probability of a sample mean score higher than 450 be smaller than, or larger than, or the same as the correct answer to part a? it is not necessary to do the detailed calculations here. Sketch a graph to illustrate your reasoning

a. Find the probability that the sample mean score is higher than 450. (I believe the answer is 0.068)
Answer #1

a)
The scores of students taking aptitude test at a particular institute, X, are normally distributed with mean 420 and standard
\sigma_{\bar{x}}=\frac{\sigma}{\sqrt{n}}=\frac{100}{\sqrt{25}}=20
a. The calculation of the probability is as shown below: z X-u P(X > 450)=1-pZ34 = 1-plz 450-420 100 25 =1-p(Z 31.5) The calc
e)
e) Consider the sample standard devotion. We know that the mean of sample standard deviation is E(52) = 100and the variance i

f)
f) We are interested in finding the value below which 5% of the sample standard deviations of scores lie. Thus we need to fin

So, we have: 2452 1002=13.85 13.85 x 1002 SE 24 = 5.770.83 = 75.97 Thus, the value, below which 5% of the sample standard dev

g)
g) Consider that instead of sample of size 25 we take a sample of size 50. It follows that the standard deviation of the samp

We can observe that the probability under the red curve is smaller than the probability under black curve

Please upvote if this helps